OpenStudy (anonymous):

can someone solve this? y= -x^2+5 together with x+4y=20

OpenStudy (lgbasallote):

solve by system of linear equation?

OpenStudy (radar):

x=0 y=5 is one answer one of the equations is not linear.

OpenStudy (radar):

\[x ^{2}+y-5=0\] The above is the first equation now multiply that by 4 getting\[4x ^{2}+4y-20=0\] Subtract the 2nd equation (in the problem) after rearranging it to look like: x+4y-20=0 after subtracting you have remaining: \[3x ^{2}=0\]\[x ^{2}=0\]x=0 then y = 5

OpenStudy (radar):

@Callum is their another root?

OpenStudy (radar):

I ask because 2nd degree equations usually have two roots.

OpenStudy (anonymous):

another way is simply substitution: \[x+4(-x^2+5) - 20 = 0\] \[x - 4x^2 +20-20 = 0\] \[4x^2 - x = 0\] \[x(4x-1) = 0\] so x = 0 or x = 1/4. So there should be a y value associated with each of these. If we let x = 0 then y = 5 by the second equation. If we let x= 1/4 then \[y=1/4(20-1/4) = 1/4(79/4) = 79/16\]. And this can be checked by putting x=1/4 into the first equation: \[y=-1/16+5 = (-1+30)/16 = 79/16\] So you have two pairs of solutions: \[(0,5)\] and \[(1/4, 79/16)\]

OpenStudy (radar):

Ah yes thanks, I hope Ilovemath sees this.

OpenStudy (anonymous):

correction: \[y=-1/16+5 = (-1+80)/16 = 79/16\]

OpenStudy (anonymous):

indeed, I think when elimination is applied to nonlinear equations there is a danger that you lose a solution...

OpenStudy (radar):

I did lol